You are eyeing an investment in a corporate bond which has a YTM of 4.00%, and a stated rate of interest of 9.00% with a maturity of 2 years. What is the price of this bond?
Answer #1

Assuming face value of bond is $1000

Coupon = 0.09 * 1000 = 90

Price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Price = 90 * [1 - 1 / (1 + 0.04)2] / 0.04 + 1000 / (1 + 0.04)2

Price = 90 * [1 - 0.924556] / 0.04 + 924.556213

Price = 90 * 1.886095 + 924.556213

Price = $1,094.30